OpenStudy (anonymous):

Solve for x. 9(x - 2) = 18

OpenStudy (gorv):

divide both side by 9

OpenStudy (anonymous):

Solve for x. 9(x - 2) = 18

OpenStudy (gorv):

divide both side by 9

OpenStudy (gorv):

\[\frac{ 9(x-2 }{ 9}=\frac{ 18 }{ 9 }\]

OpenStudy (gorv):

what u got??